OpenStudy (anonymous):

Solve the linear equation 8-5/3>9 what is the solution set

OpenStudy (lgbasallote):

can you recheck your question...i thinkk something's wrong

OpenStudy (anonymous):

My mistake its 8-5/3x>9

OpenStudy (lgbasallote):

\[8 - \frac 53 x > 9?\]

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

\[8 - \frac 53x > 9\] \[8 > 9 + \frac 53x\] \[8-9 > \frac53 x\] \[-1 > \frac 53 x\] \[-3 > 5x\] \[-\frac 35 > x\] therefore \[x < -\frac 35\]

OpenStudy (anonymous):

Thank you

OpenStudy (lgbasallote):

welcome